In another process in whic5 U undergoes neutron bombardment, the following reaction occurs. Enter the isotope symbol for the strontium (Sr) nucleus in this reaction. Express your answer as an isotope.
Answer #1

Ans. # The sum of number of protons in the “reactant” nuclei must be equal to those in product nuclei.

So,

            92 + 0 = Z + 54 + 3 x (0)

            Or, Z = 92-54 = 38

# The sum of atomic mass in the “reactant” nuclei must be equal to those in product nuclei.

            So, 235 + 1 = A + 143 + 3 x (1)

            Or, Z = 236 – 146 = 90

# Therefore isotope symbol for Sr is 90Sr38 . (please enter the both A and Z values on the left side of Sr, it’s difficult to write so in word.)